# Autoregressive models for panel data causal inference with application to state-level opioid policies

### v1

@@ Abstract
Motivated by the study of state opioid policies, we propose a novel approach that uses autoregressive models for causal effect estimation in settings with panel data and staggered treatment adoption. Specifically, we seek to estimate of the impact of key opioid-related policies by quantifying the effects of must access prescription drug monitoring programs (PDMPs), naloxone access laws (NALs), and medical marijuana laws on opioid prescribing. Existing methods, such as differences-in-differences and synthetic controls, are challenging to apply in these types of dynamic policy landscapes where multiple policies are implemented over time and sample sizes are small. Autoregressive models are an alternative strategy that have been used to estimate policy effects in similar settings, but until this paper have lacked formal justification. We outline a set of assumptions that tie these models to causal effects, and we study biases of estimates based on this approach when key causal assumptions are violated. In a set of simulation studies that mirror the structure of our application, we also show that our proposed estimators frequently outperform existing estimators. In short, we justify the use of autoregressive models to provide robust evidence on the effectiveness of four state policies in combating the opioid crisis.
Autoregressive models,
Causal inference,
@@ 1 Introduction
The U.S. continues to confront an opioid epidemic (Volkow and Collins, 2017). Between 1999 and 2019, nearly 500,000 people died from an overdose involving opioids (Centers for Disease Control and Prevention (2020a)). The number of fatal overdoses increased nearly 30% from 2019 to 2020, reflecting the spread of fentanyl and greater opioid misuse associated with the increased stress, social isolation, and job loss stemming from the COVID-19 pandemic (Centers for Disease Control and Prevention, 2020b; Patel, Walter and Li, 2021). Provisional 2024 data suggest that overdose deaths will set yet another record (Ahmad et al., 2024). In response, states continue to implement an array of policies meant to combat the opioid-crisis, producing a policy landscape that is both complex and dynamic (Davis and Carr, 2017; Davis and Lieberman, 2021; Davis and Samuels, 2020, 2021; Haffajee, Bohnert and Lagisetty, 2018; Burris et al., 2017; Smart, Haffajee and Davis, 2022; of Emergency Physicians, n.d.; Pacula and Stein, 2020; Andraka-Christou et al., 2022a, b; Frank, Humphreys and Pollack, 2021). Unfortunately, it is unclear whether any of these policies have been effective, as existing research has produced divergent effect estimates of varying credibility. Producing credible effect estimates of these policies is crucial as policymakers choose strategies to address this crisis. In this paper, we therefore quantify the effects of four policies on opioid prescriptions: must access prescription drug monitoring programs (PDMPs), two types of naloxone access laws (NALs), and medical marijuana laws.
@@
Unfortunately, obtaining credible estimates of the causal impacts of opioid-related policies is challenging. The statistical literature on estimating causal effects in panel data settings is extensive, and many existing methods are – broadly speaking – implementations of difference-in-differences (DiD) designs (Ashenfelter and Card, 1984; Angrist and Pischke, 2008) or variations of synthetic control methods (Abadie, Diamond and Hainmueller, 2010, 2015). These include extensions for effect estimation in the staggered adoption of policy setting that we consider (Callaway and Sant’Anna, 2021; Goodman-Bacon, 2021; Ben-Michael, Feller and Rothstein, 2021, 2022). Despite this, these methods are not well suited to the policy landscape seen in the opioid policy evaluation, where multiple policies are simultaneously implemented to address the opioid crisis, and disentangling these effects is important (Griffin et al., 2022). While recent approaches have tried to extend synthetic controls or DiD to multiple policies (Fricke, 2017; Agarwal, Shah and Shen, 2020; Roller and Steinberg, 2023), these are only applicable in certain situations, including when only one of the multiple treatments is of interest. Additionally, even in settings with a single policy, existing estimators have limited statistical power (Griffin et al., 2021, 2023).
@@
To address these shortcomings, Schell, Griffin and Morral (2018) and Griffin et al. (2021) proposed instead using autoregressive models for causal effect estimation. These models allow for more precise estimates of the policies effectiveness while also allowing for estimating effects of multiple policies jointly. A number of simulation studies have shown that these models perform well relative to state-of-the-art implementations of difference-in-differences and synthetic controls when evaluating policy effects (Schell, Griffin and Morral (2018)). Other studies have also shown good performance under different confounding mechanisms, the presence of co-occurring policies, and treatment effect heterogeneity (Griffin et al., 2022, 2023; Griswold et al., 2023). Unfortunately, the challenge with standard autoregressive models is that they control for lags of the dependent variable: in an analysis of policy effects, this will induce “collider bias” due to the fact that the lagged outcomes are likely affected by prior treatments (Achen, 2000; Elwert and Winship, 2014; Pearl, 2009; Nickell, 1981). So-called “debiased autoregressive models” address this problem by removing policy effects from prior outcomes in the estimation process. In this paper we formalize the assumptions behind this method that enable causal identification.
@@
Our primary methodological contribution is therefore to propose a novel causal identification result that justifies the use of debiased autoregressive models for causal effect estimation. Additionally, we study biases of estimates based on this approach when the causal assumptions are violated, particularly with respect to treatment effect heterogeneity. We find that the different estimators in this setting present a standard bias-variance trade-off and that debiased autoregressive models make a very favorable trade-off. To be precise, we show that bias is typically low for these estimators, and they can have significantly lower variance than existing estimators. This is critically important for the analysis of state opioid policy data, where both sample sizes and policy effect sizes may be small and power becomes a pressing issue. Additionally, we formalize the assumptions required to use these models to estimate effects of multiple policies simultaneously, a challenging and relevant problem in opioid policy research. While we primarily seek to demonstrate that this method is useful to quantify the impact of opioid policies, we believe that these models are useful in more general settings.
@@
Our paper proceeds as follows. In Section 2, we outline a causal framework for using debiased autoregressive models to estimate causal effects of a single policy. We also consider violations of these assumptions, particularly the potentially biasing role of effect heterogeneity. In Section 3, we relax the assumptions and show various extensions we can make to the identification result, including estimating the effects of multiple policies jointly. In Section 4, we discuss connections between our approach and other commonly used approaches, including difference-in-differences, synthetic controls, and the methods based on sequential randomization Robins, Hernan and Brumback (2000). In Section 5, we provide a set of simulation studies that further supports the use of our method for our application. In Section 6, we estimate the effects of state policies on opioid prescriptions and compare our estimates against other existing approaches. Section 7 concludes with further discussion of our methodological and applied contributions.
@@ 2.1 Simplified causal framework
In this section, we introduce a simplified version of our causal identification strategy. While facilitating exposition, this simplicity comes at the expense of strong identifying assumptions, which we relax in Section 3. Let $Y_{it}$ be an outcome for observation $i$ at time $t$, with $i=1,\dots,n$ and $t=0,\dots,T$ and let $\bar{Z}_{it}$ denote the history of some random variable $Z$ from time $t=0,...,t$. We also let $A_{it}$ denote an indicator of a policy of interest being active for observation $i$ at time $t$, and define $Y_{it}(\bar{a})$ as the potential outcome that would be observed if the treatment history $\bar{A}_{it}$ were fixed to level $\bar{a}$. Our goal is to estimate the average treatment effect at time $t$, given by $\Delta_{t}=\mathbb{E}[Y_{it}(1)-Y_{it}(0)]$, where for now we only focus on treatment at time $t$. The challenge is that we do not observe $A_{it}=1$ and $A_{it}=0$ for all individuals - in fact, we never observe both. We therefore must make assumptions tying these counterfactual quantities to the observed data distribution. We consider more complex estimands that average over multiple policy years or depend on prior treatment values in Section 3. This choice of estimand, an average treatment effect, differs from the typical estimand identified by synthetic controls and difference-in-difference estimators, which estimate a treatment effect on the treated. We further discuss differences between these approaches further in Section 4.1.
To identify this effect, we first make the following consistency assumption,
@@
Consistency states that the observed outcome under the observed policy history $\bar{A}_{it}=\bar{a}$ is equal to the potential outcome under that same policy history. This would be violated if there were multiple versions of the policy or when observed outcomes depend on other units’ policy statuses, sometimes referred to as “interference” or “spillover.” We next assume no carryover effects:
@@
For any treatment history $\bar{a}$, $Y_{it}(\bar{a})=Y_{it}(a_{t})$
No carryover effects limits the dependence of the potential outcomes on the treatment history to only the current treatment status. In other words, lagged treatments have no impact on the current outcomes. Nevertheless, this assumption places no restrictions on how treatment effects may vary over time. This assumption is common in the policy evaluation literature when using methods such as difference-in-differences and synthetic controls.
@@
In contrast to standard ignorability assumptions, we assume that randomization occurs with respect to a potentially unobserved quantity – the potential outcome under control in time $t-1$, which is unobserved for those who had already been treated. We provide a more complete discussion of this identifying assumption in comparison to other assumptions in Sections 4.1 and 4.2. We also relax this assumption to control for additional prior counterfactuals and/or covariates in Section 3. To facilitate controlling for this prior counterfactual, we next assume treatment effects are constant across units:
@@
$Y_{it}(a)=Y_{it}(0)+a\Delta_{t},\quad a=0,1.$
Constant treatment effects implies that the difference between the two counterfactuals is a fixed quantity at $t$, and it precludes any kind of random heterogeneity or heterogeneity that depends on unobservable quantities. This assumption may seem quite strong, particularly if certain units are expected to respond to treatment differently. We therefore explore violations of this assumption and how they may impact identification, along with relaxations of this assumption, in Section 2.3. Finally, we assume positivity.
@@
This requires that we have positive probability of observing either policy status conditional on the prior counterfactual under control. These core identifying assumptions allow us to write down our estimand as a function of the observed data distribution.
Theorem 1.
Under Assumptions 1-5,
$\Delta_{t}$ $=\mathbb{E}[\mathbb{E}[Y_{it}\mid A_{it}=1,Y_{i,t-1}(0)]-\mathbb{E}[Y_{it}\mid A_{it}=0,Y_{i,t-1}(0)]]$
$=\mathbb{E}\left[\mu_{1t}(Y_{i,t-1}-A_{i,t-1}\Delta_{t-1})-\mu_{0t}(Y_{i,t-1}-A_{i,t-1}\Delta_{t-1})\right],$ (1)
where $\mu_{at}(y)=E(Y_{it}\mid A_{it}=a,Y_{i,t-1}(0)=y)$. Thus, identification of the treatment effect at time $t$ requires identification of the treatment effect at time $t-1$.

### v2

@@ Abstract
Motivated by the study of state opioid policies, we propose a novel approach that uses autoregressive models for causal effect estimation in settings with panel data and staggered treatment adoption. Specifically, we seek to estimate the impact of key opioid-related policies by quantifying the effects of must access prescription drug monitoring programs (PDMPs), naloxone access laws (NALs), and medical marijuana laws on opioid prescribing. Existing methods, such as differences-in-differences and synthetic controls, are challenging to apply in these types of dynamic policy landscapes where multiple policies are implemented over time and sample sizes are small. Autoregressive models are an alternative strategy that have been used to estimate policy effects in similar settings, but until this paper have lacked formal justification. We outline a set of assumptions that tie these models to causal effects, and we study biases of estimates based on this approach when key causal assumptions are violated. In a set of simulation studies that mirror the structure of our application, we show that our proposed estimators frequently outperform existing estimators. In short, we justify the use of autoregressive models to evaluate the effectiveness of four state policies in combating the opioid crisis.
@@ 1 Introduction
The U.S. continues to confront an opioid epidemic (Volkow and Collins, 2017). Between 1999 and 2019, nearly 500,000 people died from an overdose involving opioids (Centers for Disease Control and Prevention (2020a)). The number of fatal overdoses increased nearly 30% from 2019 to 2020, reflecting the spread of fentanyl and greater opioid misuse associated with the increased stress, social isolation, and job loss stemming from the COVID-19 pandemic (Centers for Disease Control and Prevention, 2020b; Patel, Walter and Li, 2021). In response, states continue to implement an array of policies meant to combat the opioid-crisis, producing a policy landscape that is both complex and dynamic (Davis and Carr, 2017; Davis and Lieberman, 2021; Davis and Samuels, 2020, 2021; Haffajee, Bohnert and Lagisetty, 2018; Burris et al., 2017; Smart, Haffajee and Davis, 2022; of Emergency Physicians, n.d.; Pacula and Stein, 2020; Andraka-Christou et al., 2022a, b; Frank, Humphreys and Pollack, 2021). Unfortunately, it is unclear whether these policies have been effective, as existing research has produced divergent effect estimates of varying credibility. Producing credible effect estimates of these policies is crucial as policymakers choose strategies to address this crisis. In this paper, we therefore estimate the effects of four policies on opioid prescriptions: two types of naloxone access laws (NALs), must access prescription drug monitoring programs (PDMPs), and medical marijuana laws.
@@
Unfortunately, obtaining credible estimates of these effects is challenging. The statistical literature on estimating causal effects in panel data settings is extensive, and many existing methods are – broadly speaking – implementations and extensions of difference-in-differences (DiD) designs (Ashenfelter and Card, 1984; Angrist and Pischke, 2008; Callaway and Sant’Anna, 2021; Goodman-Bacon, 2021; Ben-Michael, Feller and Rothstein, 2021, 2022) and synthetic control methods (Abadie, Diamond and Hainmueller, 2010, 2015). Despite many methodological advancements in these designs, the methods are not generally well suited to the policy landscape seen in the opioid policy evaluation, where multiple policies are implemented simultaneously, and disentangling the effects of each policy is important (Griffin et al., 2022). While recent approaches have proposed extensions to synthetic controls or DiD to incorporate multiple policies (Fricke, 2017; Agarwal, Shah and Shen, 2020; Roller and Steinberg, 2023), these methods are only applicable when only one of the multiple treatments is of interest. Additionally, even in settings with a single policy, existing estimators have limited statistical power (Griffin et al., 2021, 2023).
@@
To address these shortcomings, Schell, Griffin and Morral (2018) and Griffin et al. (2021) proposed estimating effects using so-called “debiased autoregressive models,” which extend standard autoregressive models to estimate causal effects. The challenge with these standard models is that they control for lags of the dependent variable: in an analysis of policy effects, this will induce “collider bias” due to the fact that the lagged outcomes are likely affected by prior treatments (Achen, 2000; Elwert and Winship, 2014; Pearl, 2009; Nickell, 1981). At a high level, debiased models address this problem by removing policy effects from prior outcomes in the estimation process. While lacking formal justification, a number of simulation studies showed that these models performed well relative to state-of-the-art implementations of difference-in-differences and synthetic controls when evaluating policy effects (Schell, Griffin and Morral (2018)). Other studies demonstrated good performance under different confounding mechanisms, the presence of co-occurring policies, and treatment effect heterogeneity (Griffin et al., 2022, 2023; Griswold et al., 2023).
@@
Our primary methodological contribution is therefore to propose a novel causal identification result that justifies the use of debiased autoregressive models for estimating causal effects. Additionally, we study biases of estimates based on this approach when the causal assumptions are violated, particularly with respect to treatment effect heterogeneity. We find that the different estimators in this setting present a standard bias-variance trade-off and that debiased autoregressive models make a very favorable trade-off. To be precise, we show that bias is typically low for these estimators, and they can have significantly lower variance than existing estimators. This is critically important for the analysis of state opioid policy data, where both sample sizes and policy effect sizes may be small. Additionally, we formalize the assumptions required to use these models to estimate effects of multiple policies simultaneously, a challenging and relevant problem in opioid policy research. While we primarily seek to demonstrate that this method is useful to quantify the impact of opioid policies, these models are likely useful in both other policy evaluations and more general settings.
@@
Our paper proceeds as follows. In Section 2, we introduce the state-level opioid data and the corresponding data structure for our application. In Section 3, we outline a simplified causal framework that motivates using debiased autoregressive models to estimate causal effects of a single policy. In Section 4, we discuss several extensions that relax several of the assumptions from Section 3, including estimating the effects of multiple policies jointly, and a discussion of the potential for bias in the presence of effect heterogeneity. In Section 5, we discuss connections between our approach and other commonly used approaches, including difference-in-differences, synthetic controls, and methods based on sequential randomization (Robins, Hernan and Brumback, 2000). In Section 6, we provide a set of simulation studies that further supports the use of our method for our application and highlights when this approach may perform well. In Section 7, we estimate the effects of state policies on opioid prescriptions and compare our estimates against other existing approaches. Section 8 concludes with further discussion of our methodological and applied contributions.
2 State-level opioid policy data
Our work is motivated by the need to assess the impact of four state-level policies, using data on the number of opioid prescriptions per capita in each state over the years 2006-2017. In this section, we provide background on the data and its structure.
2.1 Policies, outcomes, and covariates
We obtained our policy measures from the OPTIC-Vetted Policy Data Sets from the RAND-USC Opioid Policy Tools and Information Center (OPTIC) (RAND-USC Schaeffer Opioid Policy Tools and Information Center, 2024a, b, c). We study four policies that have been examined in prior research but are rarely considered jointly: two types of naloxone access laws (NALs), must access prescription drug monitoring programs, and medical marijuana laws.
Naloxone is a potential life saving treatment for opioid overdoses that works by reversing the effects of opioids (Cawley and Dragone, 2023; Chamberlain and Klein, 1994). However, naloxone may be difficult to obtain due to stigma or lack of access to traditional healthcare systems (Green et al., 2020; Smart and Grant, 2021). To address these barriers, every state has adopted some type of NAL by 2018 (Smart, Pardo and Davis, 2021). There are two broad classes of NALs (Smart et al., 2023): those which allow distribution through a standing or protocol order or those which allow distribution through a pharmacist prescriptive authority (Davis and Carr, 2015, 2017). We therefore consider each class as a separate policy. PDMPs are electronic databases that track prescribed and dispensed controlled substances within a state (Schuler et al., 2020). By giving healthcare providers information about a patient’s prescription drug history, PDMPs may help identify patients who are at risk for opioid misuse or overdose, potentially increasing early intervention and support (Horwitz et al., 2018). Based on the literature on PDMPs, state laws that specify that prescribers must access the PDMP before writing a prescription to a new/existing patient tends to allow for more effective forms of PDMPs (Pacula and Stein, 2020; RAND-USC Schaeffer Opioid Policy Tools and Information Center, 2024b). Thus, we define our PDMP laws as those which have this must access component. Finally, medical marijuana laws legalize marijuana for medical purposes (Pacula et al., 2015). These laws may reduce opioid use and opioid-related harms by providing an alternative treatment option for chronic pain and other conditions that are commonly treated with opioids (Powell, Pacula and Jacobson, 2018). For a summary of the prevalance of these laws over time, see Figure 1 below. [^1] For all four policies, our policy measure is the fraction of the year each state has such a law in place.
Figure 1: Policy wheels showing the timeline of policy adoption
We examined the impact of these policies on annual CDC opioid prescribing rates per 100,000 people from 2006-2017. In contrast to the policy measures, we construct the outcomes from the QuintilesIMS Transactional Data Warehouse (TDW). These data represent approximately 59,000 retail (non-hospital) pharmacies, which dispense nearly 88% of all retail prescriptions in the U.S. We identify opioid prescriptions using the National Drug Code. We do not include cough and cold formulations, which contain opioids, or buprenorphine, an opioid partial agonist used for treatment of opioid use disorder and pain. To convert the prescription counts into rates, we use annual resident population estimates obtained from the Population Estimates Program, U.S. Census Bureau. An illustration of opioid prescription rates over time by state can be found in Appendix E.
Finally, we include seven time-varying covariates in our models. These are: annual unemployment rates from the St. Louis Federal Reserve’s FRED Economic Data system, median income and poverty rate derived from the U.S. Census Bureau’s 1-year American Community Survey (ACS), and percentages of non-Hispanic white individuals, females, individuals under 20 years old, and those over 64 years old, derived from the U.S. Census Intercensal (2000–2010) and Postcensal (2011–2022) estimates. These covariates were selected to capture a range of sociodemographic factors that might confound the effects of opioid policies on the outcomes.
2.2 Data structure and summary
Our data include 51 units (50 states plus Washington D.C.) observed over 12 years at yearly intervals. An illustration of the structure of our data for one of the policy variables is given in Table 1, which illustrates that the policy variables follow a staggered adoption regime, where states adopt policies at different times, and policy adoption does not reverse during the course of our study. For more information on all of the variables in our data, Table 2 displays summary statistics for each variable in both the first and last year of our study. Opioid prescription rates drop over the course of our study. Additionally, at the beginning of the study, only medical marijuana laws were enacted among all four policies, whereas by the end many states had enacted these policies. Covariate values follow expected trends over time, such as increasing income over time.
2006 2007 2008 2009 2010 2016 2017
State 1
State 2
State 50
State 51
Table 1: Hypothetical illustration of data structure for a single policy. Checks indicate a year in which the policy was enacted. Note there are 51 states as we include Washington D.C as a unit of analysis.
Year 2006 Year 2017
Covariates Median income (\$) 24154 (3239) 30407 (4469)
Unemployment rate (%) 4.48 (1.08) 4.16 (0.93)
Percentage White (%) 72.43 (16.10) 68.26 (16.16)
Percentage female (%) 50.74 (0.79) 50.59 (0.84)
Percentage less than 20 (%) 27.33 (1.90) 25.15 (2.07)
Percentage above 64 (%) 12.65 (1.67) 15.96 (1.91)
Percentage in poverty (%) 12.67 (3.07) 12.78 (2.82)
Policies Naloxone standing (%) 0 (0) 85.29 (32.13)
Naloxone prescriptive (%) 0 (0) 15.69 (34.55)
PDMP must access laws (%) 0 (0) 33.33 (47.61)
Medical marijuana (%) 23.53 (42.84) 56.21 (49.66)
Outcome Prescription rate (per 100,000) 76.71 (21.69) 62.59 (17.91)
Table 2: Mean and standard deviation of variables in our study in 2006 and 2017
@@ 3 Methods
We now motivate our methodological approach, debiased autoregressive models, and first introduce a simplified version of our causal identification strategy, which suggests a simple parametric modeling approach to estimation. While facilitating exposition, this simplicity comes at the expense of strong identifying assumptions and restrictive model parameterizations. We defer outlining relaxations of these assumptions to Section 4, though we rely on these relaxations for our analysis of state-level opioid data in Section 7. We begin by formalizing the causal estimands that we wish to estimate, and then outline the set of simplified assumptions that allow us to identify these estimands in terms of the observed data distribution.
@@ 3.1 Simplified causal framework
3.1.1 Estimands
We consider a setting where we observe $Y_{it}$, an outcome of the observation $i$ at time $t$, with $i=1,\dots,n$ and $t=0,\dots,T$. We let $\bar{Z}_{it}$ denote the history of some random variable $Z$ from time $t=0,\dots,t$. We also let $A_{it}$ denote an indicator of a single policy of interest being active for observation $i$ at time $t$, and define $Y_{it}(\bar{a})$ as the potential outcome that would be observed if the treatment history $\bar{A}_{it}$ was fixed at the level $\bar{a}$. Finally, we let $T_{0i}+1$ denote the time that unit $i$ first adopts policy $A_{it}$. In contrast to our real data structure, we do not yet consider other policies or covariates. Our goal is to estimate the average treatment effect across all study time periods, given by
$\theta=\mathbb{E}[Y_{it}(1)-Y_{it}(0)].$
This estimand makes no restrictions on heterogeneity over time or across units, though averages over both. Throughout this section and Section 4, we focus on this population level estimand as it easily allows us to study identification and biases under violations of our assumptions. At times in the manuscript, however, we instead consider a sample-level average treatment effect on the treated (SATT),
$\theta^{\mathcal{S}}_{m}=\frac{1}{n_{t}}\sum_{i\in\mathcal{N}_{t}}\sum_{m=1}^{M}\bigg\{Y_{i,T_{0i}+m}(1)-Y_{i,T_{0i}+m}(0)\bigg\},$
where $\mathcal{N}_{t}$ contains the indices of treated units, and $n_{t}$ is the number of treated units. This estimand reflects the average treatment effect on the treated units in the first $M$ time periods after policy adoption. Our proposed framework allows for estimation of either quantity, though we leave details on performing inference on $\theta^{\mathcal{S}}_{m}$ to Appendix C. Many popular methods, such as difference-in-differences and synthetic controls, target $\theta^{\mathcal{S}}_{m}$ in the staggered adoption setting. We therefore consider this quantity in order to compare our approach to other existing approaches in simulations and for the opioid policy application. However, we also consider more complex estimands that examine multiple policies simultaneously or depend on prior treatment values in Section 4.
A major challenge when estimating these causal effects is that we never observe the outcomes when $A_{it}=1$ and $A_{it}=0$ for any individuals. We therefore must make assumptions tying both potential outcomes $Y_{it}(1)$ and $Y_{it}(0)$ to the observed data distribution. We discuss a set of assumptions sufficient to identify $\theta$ in terms of the observed data distribution below.
3.1.2 Assumptions
We again first introduce a simple set of assumptions that are sufficient to identify $\theta$ in terms of the observed data distribution, and we defer relaxing these assumptions to Section 4. We first make the consistency assumption,
@@ 3.1.2 Assumptions
Consistency states that the observed outcome under the observed policy history $\bar{A}_{it}=\bar{a}$ is equal to the potential outcome under that same policy history. This would be violated if there were multiple versions of the policy or when observed outcomes depend on other units’ policies, sometimes referred to as “interference” or “spillover effects.”
We next assume no carryover effects,
@@
For any treatment history $\bar{a}$, $Y_{it}(\bar{a})=Y_{it}(a_{t})$.
No carryover effects limits the dependence of the potential outcomes on the treatment history to only the current treatment status. In other words, we assume that prior treatments have no impact on the current outcomes. While strong, this assumption places no restrictions on how treatment effects may vary over time, and is common in the policy evaluation literature.
@@
In contrast to standard ignorability assumptions, we assume that randomization occurs with respect to a potentially unobserved quantity, the prior control counterfactual, which is unobserved for those who had already been treated. We provide a more complete discussion of this identifying assumption in comparison to other assumptions in Sections 5.1 and 5.2.
We next assume treatment effects are constant across units:
@@
$Y_{it}(a)=Y_{it}(0)+a\theta,\quad a=0,1.$
Constant treatment effects implies that the difference between the two counterfactuals is a fixed quantity across units and time (i.e. $Y_{it}(1)-Y_{it}(0)=\theta$) and it precludes any kind of random heterogeneity. This assumption is especially strong, particularly if certain units are expected to respond to treatment differently.
We also assume positivity,
@@
This requires that we have positive probability of observing either policy status conditional on the prior counterfactual under control.
Finally, we assume that the expected potential outcomes follow some known parametric model.
Assumption 6 (Known parametric model).
Assume that for each unit $i$ and time period $t$
$Y_{it}(0)=\alpha_{t}+\beta_{t}^{\top}\phi(Y_{i,t-1}(0))+\epsilon_{it},$ (1)
for some parameter vector $(\alpha_{1},\beta_{1}^{\top},\dots,\alpha_{T},\beta_{T}^{\top})^{\top}\in\mathbb{R}^{T(q+1)}$ and some known function $\phi(x):\mathbb{R}\to\mathbb{R}^{q}$. Let $Z_{it}=[1,A_{it},\phi^{\top}(Y_{it-1}(0))]$ and assume that $\mathbb{E}[Z_{it}^{\top}Z_{it}]$ has full column rank.
Proposition 1.
Under assumptions 1-6, $\theta$ is identified in terms of the observed data distribution as the minimizer of $\mathbb{E}[(Y_{it}-\theta A_{it}-\alpha_{t}-\beta_{t}^{\top}\phi(Y_{i,t-1}-\theta A_{i,t-1}))^{2}]$.
There is no general non-parametric identification result in this setting, so our proposed identification strategy relies on parametric models for identification. While the parametric assumptions may seem particularly strong, in practice one may choose a flexible set of functions $\phi$ to model the outcomes. Moreover, in many common scenarios $\theta$ may be identified non-parametrically. For example, there is often a time period $t^{\star}$ in which no units are treated, so that $Y_{i,t^{\star}}(0)=Y_{i,t^{\star}}$ for all units. In this case, Proposition 2 shows that non-parametric identification can be obtained.
Proposition 2.
Assume (1)-(5), and that there exists some time period $t^{\star}$ where $A_{it^{\star}}=0$ for all units ii. Then,
$\theta=\mathbb{E}[\mathbb{E}[Y_{i,t^{\star}+1}\mid A_{i,t^{\star}+1}=1,Y_{it^{\star}}]-\mathbb{E}[Y_{i,t^{\star}+1}\mid A_{i,t^{\star}+1}=0,Y_{it^{\star}}]].$ (2)
Equation (2) in Proposition 2 shows that $\theta$ can be identified in the data by modeling the conditional expectation function using outcomes from the $t^{\star}+1$ time period. Additionally, in some cases we may want to assume that $\theta$ varies over time – i.e. we may assume that equation (2) only identifies $\theta_{t^{\star}+1}$. In this case, we may recursively identify $\theta_{t}$ for any time period $t>t^{\star}+1$ using
$\theta_{t}=\mathbb{E}[\mathbb{E}[Y_{it}\mid A_{it}=1,Y_{i,t-1}-A_{i,t-1}\theta_{t-1}]-\mathbb{E}[Y_{it}\mid A_{it}=0,Y_{i,t-1}-A_{i,t-1}\theta_{t-1}]],$
suggesting a corresponding estimation procedure that iteratively estimates each $\theta_{t}$. Note that this non-parametric identification strategy only uses data from a single time period to estimate the effect at that time, which leads to a reduction in efficiency. Because of this, we focus on our parametric modeling approach, which borrows information across time periods by assuming a constant treatment effect across time, or by assuming it can be modeled as a parametric function of time as we will see in subsequent sections.